# Quasi-Solid and Supersolid from Quasiperiodic Long-Range Interactions

### v1

@@ Abstract
We investigate hard-core bosons in one dimension with quasiperiodic long-range interactions, defined by $V_{ij}=V_{0}\cos(\pi\alpha i)\cos(\pi\alpha j)$ with the inverse golden ratio $\alpha=(\sqrt{5}-1)/2$. Large-scale quantum Monte Carlo simulations reveal incompressible density plateaus at incommensurate fillings tied to Fibonacci ratios. These plateaus feature emergent nonuniform density profiles and robust long-range correlations, as captured by the structure factor. Depending on filling and interaction strength, the system realizes either a quasi-solid phase with suppressed superfluidity, a quasi-supersolid phase where density order coexists with finite superfluid density, or a superfluid phase. Our results demonstrate that purely interaction-induced quasiperiodicity, without external potential or disorder, can stabilize novel quantum phases that simultaneously break translational symmetry and sustain quantum coherence.
Introduction– Quasiperiodic systems provide a unique platform to study quantum phases at the interface of order and disorder. Unlike random disorder, quasiperiodic patterns are deterministic yet aperiodic, producing incommensurate spatial structures without introducing randomness. Such quasiperiodic structures have been extensively explored in low-dimensional quantum systems, where modulations based on Aubry–André [1, 2, 3, 4, 5, 6] or Fibonacci sequences [7, 8, 9] give rise to striking phenomena including single-particle spectral criticality [10, 11, 12], mobility edges [13, 14, 15, 16], topological band structures [17, 18, 19], and many-body localization [20, 21, 7, 9, 22]. These studies, however, have almost exclusively relied on imposing quasiperiodicity through external on-site potentials or lattice modulations.
In contrast to the extensively studied effects of quasiperiodic potential, relatively little attention has been paid to systems where the interparticle interactions themselves exhibit deterministic quasiperiodicity. This alternative pathway to unconventional quantum phases circumvents the need for single-particle localization or extrinsic disorder, instead harnessing spatially modulated, incommensurate couplings to induce emergent collective behavior.
Recent theoretical work has proposed that quasiperiodic long-range interactions can be engineered using multi-mode cavity interference, enabling real-space interaction profiles that approximate Fibonacci-like modulations even in the absence of external quasiperiodic potentials [23]. While that study focuses on localization and entanglement dynamics in a fermionic setting, our work demonstrates that such interaction structures can also stabilize exotic ordered phases, including quasi-solids and quasi-supersolids, in bosonic systems at equilibrium.
At the same time, rapid advances in programmable quantum simulators have brought the physical realization of interaction-engineered models within reach. Notably, Rydberg-atom arrays with tunable blockade radii [24, 25] and cavity-mediated atomic ensembles with controllable mode structures [26, 27] now provide site-resolved access to both the range and spatial profile of interactions. These capabilities enable the deterministic implementation of aperiodic couplings, either through tailored atomic geometries or through Fourier-space engineering of the interaction kernel. In particular, the realization of Fibonacci-modulated interactions, long considered theoretically intriguing, now appears experimentally feasible, significantly enhancing the relevance and testability of models with quasiperiodic long-range interactions.
In this work, we investigate a one-dimensional hard-core Bose–Hubbard model with quasiperiodic long-range interactions. Crucially, our model contains neither on-site potential disorder nor on-site repulsive terms, thereby isolating the role of quasiperiodic interactions alone. Using large-scale quantum Monte Carlo simulations with the worm algorithm [28, 29], we demonstrate that this deterministic interaction pattern stabilizes two striking states: (i) a quasi-solid phase, marked by incompressible density plateaus at Fibonacci fillings and sharp incommensurate peaks in the structure factor, and (ii) a quasi-supersolid phase, where such density modulations coexist with finite superfluid density. These results establish interaction-driven quasiperiodicity as a new paradigm for engineering exotic ordered states beyond conventional supersolids.
Our findings not only broaden the theoretical landscape of quasiperiodic systems but also directly connect to ongoing experimental efforts in engineered long-range interactions. By demonstrating how deterministic aperiodicity in the interaction sector alone can stabilize incompressible and supersolid-like phases, we provide a roadmap for realizing and probing such exotic states in next-generation quantum simulators.
@@
To identify different phases, we compute the following observables. The average filling is defined as $\langle n\rangle=\frac{1}{L}\sum_{x}\langle n_{x}\rangle$, and incompressible plateaus in $\langle n\rangle(\mu/V)$ indicate gapped states. The compressibility is $\kappa=\frac{\beta}{L}(\langle N^{2}\rangle)-\langle N\rangle^{2}$, with $N$ total number of particles of the system, estimated via density fluctuations, with vanishing $\kappa$ marking insulating behavior. The superfluid density is $\rho_{s}=\langle W^{2}\rangle/(2\beta t)$ [30], where $W$ is the winding number of worldlines. Here, $\rho_{s}>0$ implies global phase coherence.
To characterize spatial ordering we compute the imaginary-time–averaged connected density–density correlator $C(x)=\frac{1}{L}\sum_{i}\big(\langle n_{i}n_{i+x}\rangle-\langle n_{i}\rangle\langle n_{i+x}\rangle\big)-\langle n\rangle$. Its Fourier transform, the static structure factor, $S(q)=\frac{1}{L}\sum_{x}e^{iqx}C(x)$, provides momentum-resolved information on density modulations. In periodic systems $S(q)$ exhibits Bragg peaks at commensurate momenta, while in quasiperiodic systems sharp incommensurate peaks signal quasi–long-range order. In our model these peaks appear at golden–ratio–related wavevectors $q^{*}$, offering a clear signature of quasi-solid or quasi-supersolid phases stabilized by interaction-driven quasiperiodicity.
@@
Figure 1: Ground-state phase diagram of the system in the $\mu/V$$t/V$ plane for $L=233$. Two incompressible quasi-solid (qS) lobes emerge near fillings $\langle n\rangle=0.618\approx 1/\alpha$ and $\langle n\rangle=0.382\approx 1-1/\alpha$, where $\alpha=(1+\sqrt{5})/2$ is the golden ratio. These phases exhibit vanishing compressibility and superfluid density, along with sharp incommensurate peaks in $S(q)$, indicating Fibonacci-induced spatial order. The light blue region marks the quasi-supersolid (qSS) regime, where incommensurate density modulations coexist with finite superfluid density.
Phase diagram– Figure 1 summarizes the ground-state phase diagram of a one-dimensional hard-core Bose system with Fibonacci-modulated quasiperiodic long-range interactions. For system size $L=233$, we map out the phases in the $\mu/V$$t/V$ plane and identify two insulating lobes near fillings $\langle n\rangle\approx 1/\alpha$ and $1-1/\alpha$, where $\alpha$ is the golden ratio. These lobes exhibit zero compressibility ($\kappa=0$), vanishing superfluid density ($\rho_{s}=0$), and sharp incommensurate Bragg peaks in S(q)S(q)—hallmarks of quasi-solid phases stabilized solely by deterministic incommensurate interactions.
@@
Figure 2: Average filling $\langle n\rangle$ (green circles), superfluid density $\rho_{s}$ (red squares), and structure factor at the dominant incommensurate peak $S(q^{*})$ (purple upward trangules) as functions of chemical potential $\mu/V$ for $L=233$ at $t/V=0.5$. Two prominent incompressible plateaus appear at irrational fillings $\langle n\rangle=0.618$ and $\langle n\rangle=0.382$, corresponding to the golden-ratio fractions $1/\alpha$ and $1-1/\alpha$ with $\alpha=(1+\sqrt{5})/2$. These densities are naturally commensurate with the underlying Fibonacci-modulated interaction pattern and host quasi-solid order, characterized by vanishing $\rho_{s}$ and enhanced $S(q^{*})$. Between these plateaus, finite $\rho_{s}$ coexists with nonzero $S(q^{*})$, signaling the emergence of a quasi-supersolid phase stabilized purely by quasiperiodic long-range interactions. Error bars are within symbols if not seen in the figure.
@@
Figure 3: (a) Site-resolved average density profile $\langle n_{x}\rangle-\langle n\rangle$ along a one-dimensional chain of length $L=233$, with total particle number $N=144$, corresponding to filling $\langle n\rangle=0.618\approx 1/\alpha$, where $\alpha=(1+\sqrt{5})/2$ is the golden ratio at $t/V=1/3$. The quasiperiodic long-range interaction follows a Fibonacci modulation pattern, leading to a quasi-solid phase characterized by incommensurate density modulations, vanishing superfluid density, and zero compressibility. (b) Zoom-in of sites 1–54 of the same system. The circle radii are proportional to the local density at each site, emphasizing the incommensurate spatial modulation.
@@
Figure 4: (a) Density–density correlation function $C(x)$ at $t/V=1/3$ for three representative fillings on a system of size $L=233$: $\langle n\rangle=0.618$ (red circles), $\langle n\rangle=0.694$ (blue squares), and $\langle n\rangle=0.780$ (green upward triangles). Red data correspond to the inverse golden ratio $\langle n\rangle=0.618\approx\alpha^{-1}$, at which quasiperiodic modulation is most pronounced. Fitting curves are superimposed on the data. (b) Corresponding structure factor $S(q)$ for each filling, displaying prominent peaks at incommensurate wavevectors $q_{1}=2.398$ and $q_{2}=3.882$ that reflect the dominant Fourier components of the Fibonacci pattern. Inset: finite-size scaling of the dominant peak height $S(q^{*})$ at $q^{*}=2.398$ for $\langle n\rangle=0.618$ with the fitting function as $S(q^{*})(L)=0.126991+14.4886\,\left(\frac{1}{L}\right)^{1.45515}$, which confirms the persistence of quasi-long-range order in the thermodynamic limit.
@@
To further characterize the nature of the emergent quasi-solid and quasi-supersolid phases, we analyze the time averaged connected density–density correlation function $C(x)=\frac{1}{L}\sum_{i}(\langle n_{i}n_{i+x}\rangle-\langle n_{i}\rangle\langle n_{i+x}\rangle)-\langle n\rangle$ and its Fourier transform $S(q)$ at $t/V=1/3$ for system size $L=233$. Figure 4(a) shows $C(x)$ at three representative fillings: $\langle n\rangle=0.618$, $0.694$, and $0.780$. At $\langle n\rangle=0.618\approx\alpha^{-1}$, $C(x)$ displays coherent quasiperiodic oscillations over long distances, in stark contrast to the more damped behavior at nearby fillings. We fit $C(x)$ using the function $C(x)=A\cos(qx)e^{-x/\xi}+C_{0}$ for each case. For $\langle n\rangle=0.618$, the best-fit parameters are $A=0.0837$, $q=2.4$, $\xi\approx 185001$, and $C_{0}\approx-2.34\times 10^{-7}$; for $\langle n\rangle=0.694$, the fitting yields $A=0.0334$, $q=2.40001$, $\xi\approx 31615.1$, and $C_{0}\approx-5.34\times 10^{-7}$; while for $\langle n\rangle=0.780$, we obtain $A=0.0122$, $q=2.40001$, $\xi\approx 4962.7$, and $C_{0}\approx-1.21\times 10^{-6}$. The gradual reduction in both the amplitude $A$ and correlation length $\xi$ as the filling deviates from the Fibonacci value does not imply the disappearance of ordered phases. Rather, it reflects a smooth transition from a strongly correlated quasi-solid state at $\langle n\rangle=0.618$—where density modulations dominate and $\rho_{s}=0$—to a quasi-supersolid regime at nearby incommensurate fillings. In this quasi-supersolid regime, the modulated density pattern persists while phase coherence emerges, as evidenced by finite values of superfluid density $\rho_{s}=0.297\pm 0.0008$, compressibility $\kappa=0.323\pm 0.007$ at $\langle n\rangle=0.694$, and $\rho_{s}=0.247\pm 0.0003$, $\kappa=0.384\pm 0.009$ at $\langle n\rangle=0.780$. This evolution highlights the robustness of quasiperiodic interaction–induced ordering: while the strength of density modulation weakens away from the Fibonacci resonance, the system continuously retains quasi-supersolid phase at $t/V=1/3$.
@@
For a system of length $L=F_{k}$, where $F_{k}$ denotes the kk‑th Fibonacci number, the leading Fourier components of the quasiperiodic modulation occur at $q^{*}=2\pi\frac{F_{k-2}}{F_{k}}$ and its mirror $2\pi-q^{*}$. These characteristic momenta arise from the Fourier transform of the Fibonacci characteristic sequence and are a hallmark of its singular‑continuous diffraction spectrum [31, 32, 33]. In our case with $L=233=F_{13}$, the dominant peak appears at $q^{*}=2.398$, corresponding to $2\pi\times 89/233$, confirming that the observed density pattern is intrinsically locked to the golden ratio $\alpha=(1+\sqrt{5})/2$ through its rational approximants $F_{k-2}/F_{k}$. The self‑similar nature of the Fibonacci sequence therefore governs the emergence of quasi‑long‑range spatial order. In addition to these principal peaks, two weaker features appear at $q_{3}=1.483$ and $q_{4}=4.800$. These originate from higher‑order Fourier components of the singular‑continuous spectrum — specifically from secondary harmonics associated with other Fibonacci ratios such as $F_{k-3}/F_{k}$, and reflect the hierarchical, self‑similar structure of the underlying quasiperiodic modulation.
Upon doping away from $\langle n\rangle=1/\alpha$, the Bragg peaks gradually decreases while superfluid density remains finite, signaling a transition to a quasi-supersolid regime. This demonstrates that quasiperiodic long-range interactions can stabilize both incompressible quasi-solids and nearby quasi-supersolid phases with simultaneous density order and phase coherence.
Discussion and Outlook– Our results demonstrate that deterministic quasiperiodic long-range interactions, even in the absence of external potentials, disorder, or soft-core effects, are sufficient to stabilize exotic many-body phases—specifically, quasi-solids and quasi-supersolids. These phases emerge from the incommensurate structure embedded in the interaction profile and are characterized by the coexistence of spatial density modulation and long-range superfluid coherence.
@@
“Quasi-Solid and Supersolid from Quasiperiodic Long-Range Interactions”
@@
with total particle number $N=\sum_{x}\langle n_{x}\rangle$. Vanishing $\kappa$ signals incompressibility.
@@
In particular, for $L=233=F_{13}$, the dominant peaks appear at $q^{*}=2\pi\cdot 89/233=2.398$ and its mirror $2\pi-q^{*}=3.882$, consistent with the main features observed in $S(q)$. This confirms that the emergent density order is intrinsically locked to the golden ratio $\alpha=(1+\sqrt{5})/2$ via its rational approximants $F_{k-2}/F_{k}$.
In addition to these dominant peaks, we also observe subleading features at $q=1.483$ and $4.800$, which correspond to higher-order harmonics at $2\pi\cdot F_{k-3}/F_{k}=2\pi\cdot 55/233$ and its mirror. Together, the presence of both dominant and subleading peaks confirms that the system exhibits quasiperiodic density order governed by multiple competing incommensurate wavevectors.
@@
Figure 5: (a) Average filling $\langle n\rangle$, superfluid density $\rho_{s}$, and structure factor $S(q^{*})$ as functions of the chemical potential $\mu/V$ for $t/V=2$ at system size $L=233$. The system is predominantly superfluid, exhibiting finite $\rho_{s}$, while $S(q^{*})$ remains nearly zero across most of the chemical potential range. A weak enhancement of $S(q^{*})$ ($\max S(q^{*})\approx 0.035$) occurs near the golden-ratio fillings $\langle n\rangle=0.382$ and $\langle n\rangle=0.618$, indicating the quasiperiodic order. (b) Corresponding data for $t/V=1/3$. In this regime, well-defined incompressible plateaus appear at $\langle n\rangle=0.382$ and $\langle n\rangle=0.618$, accompanied by enhanced $S(q^{*})$, marking the emergence of stable quasi-solid phases. Between the plateaus, $\rho_{s}$ remains finite while $S(q^{*})$ stays nonzero, signifying a quasi-supersolid region where phase coherence coexists with incommensurate density modulation.
@@
Weak Interaction Regime (t/V=2t/V=2)– Figure S1(a) shows the behavior of $\langle n\rangle$, $\rho_{s}$, and $S(q^{*})$ as functions of $\mu/V$ in the weak-interaction regime. The system is predominantly superfluid: $\rho_{s}$ is finite, and $S(q^{*})$ stays very small over most of the $\mu/V$ range. However, near fillings $\langle n\rangle=0.618\approx\alpha^{-1}$ and $0.382=1-\alpha^{-1}$, $S(q^{*})$ develops a shallow maximum (peaking at $\sim 0.035$) and $\langle n\rangle(\mu/V)$ exhibits narrow plateaus. According to our conservative criterion ($S(q^{*})>0.03$), these windows mark the onset of quasi-solid correlations and incipient incompressibility, identifying them as quasi solid regions. Outside these plateaus, $S(q^{*})$ quickly falls below 0.03, indicating that the system remains a compressible superfluid.

### v2

# Interaction-Induced Quasicrystalline Order: Emergence of Quasi-Solid and Quasi-Supersolid Phases

@@ Abstract
Deterministic quasiperiodicity in quantum systems has long been associated with localization, criticality, or glassy behavior, and has therefore been believed to suppress long-range order rather than stabilize it. Here we demonstrate the opposite: quasiperiodicity in interactions—without any quasiperiodic potential, disorder, or geometric modulation—can generate coherent, ordered quantum phases. We study hard-core bosons in one dimension with quasiperiodic long-range interactions, $V_{ij}=V_{0}\cos(\pi\alpha i)\cos(\pi\alpha j)$, where $\alpha=(\sqrt{5}-1)/2$ is the inverse golden ratio. Using large-scale path-integral quantum Monte Carlo simulations, we uncover thermodynamically stable incompressible plateaus at irrational densities tied to Fibonacci ratios. These plateaus exhibit sharp incommensurate Bragg peaks, signaling an emergent quasi-solid with long-range quasicrystalline density order. More strikingly, at nearby fillings and interaction strengths, we identify a quasi-supersolid phase that supports both Fibonacci density ordering and finite superfluid density—demonstrating that interaction-induced quasiperiodicity can stabilize supersolid coherence. Our results establish a new mechanism for realizing ordered quasicrystalline quantum matter, and provide realistic guidance for implementation in Rydberg atom arrays, multimode cavity-QED systems, and trapped-ion quantum simulators.
Introduction– Quasiperiodic quantum systems sit at the intriguing interface between order and disorder: deterministic yet aperiodic, neither random nor periodic, they challenge the conventional paradigms of quantum matter. Over the past four decades, quasiperiodicity has become synonymous with localization [1, 2], criticality [3, 4, 5], and Bose-glass phases [6, 7, 8, 9, 10, 11], where long-range coherence is suppressed rather than stabilized. From Aubry–André potentials and bichromatic lattices to Fibonacci sequences and quasicrystalline geometries, studies have consistently concluded that quasiperiodicity destroys order—not creates it [1, 2, 6, 7, 8, 9, 10, 11, 12, 3, 4, 13, 14, 15, 16, 17, 18].
But is this always true? Can quasiperiodicity, under the right conditions, generate—rather than suppress—quantum order? This core question has remained essentially unexplored because quasiperiodicity has almost exclusively been implemented through external on-site potentials or lattice geometry, rather than through the interactions themselves. Potential-driven quasiperiodicity inherently competes with coherence, favoring localization, multifractality, or glassy behavior, even in strongly interacting regimes. What has been largely overlooked is a fundamentally different mechanism: deterministic quasiperiodicity embedded directly in the interaction kernel, instead of the on-site potential.
Recent theoretical work has proposed that quasiperiodic long-range interactions can be engineered, for example, using multimode cavity interference, yielding real-space interaction profiles that approximate Fibonacci-like modulations even in the absence of external quasiperiodic potentials [19]. At the same time, rapid advances in programmable quantum simulators—including Rydberg-atom arrays with tunable blockade radii [20, 21], cavity-mediated atomic ensembles with controllable mode structures [22, 23], and trapped-ion platforms with engineered power-law interactions [24]—have brought the physical realization of interaction-engineered models within reach. These developments naturally raise a profound and previously unresolved question: Can deterministic quasiperiodicity in the interaction sector alone stabilize ordered quantum phases—including solids or supersolids—that have long been believed incompatible with quasiperiodic systems?
In this work, we answer this question in the affirmative. We investigate a one-dimensional hard-core Bose–Hubbard model with purely interaction-induced quasiperiodicity, defined by $V_{ij}=V_{0}\cos(\pi\alpha i)\cos(\pi\alpha j),$ where $\alpha=(\sqrt{5}-1)/2$ is the inverse golden ratio. Crucially, our model contains neither on-site potential disorder nor quasiperiodic on-site modulation, thereby isolating the role of quasiperiodic long-range interactions alone.
Using large-scale path-integral quantum Monte Carlo simulations with the worm algorithm [25, 26], we uncover two thermodynamically stable quantum phases that, to the best of our knowledge, have not been observed in any previous quasiperiodic system: (i) a quasi-solid phase, marked by incompressible density plateaus at Fibonacci fillings and sharp incommensurate peaks in the structure factor, and (ii) a quasi-supersolid phase, where such quasicrystalline density modulations coexist with finite superfluid density. These results demonstrate that quasiperiodicity in the interaction kernel—rather than suppressing coherence—can drive the emergence of ordered and even supersolid quantum matter, overturning the conventional belief that quasiperiodicity inevitably leads to glassiness or localization.
Moreover, because the required quasiperiodic interaction patterns can be engineered using Rydberg-atom arrays, multimode cavity-QED systems, and trapped-ion simulators, our findings provide a realistic roadmap for realizing and probing quasicrystalline solids and supersolids in next-generation quantum simulators. In this sense, interaction-driven quasiperiodicity defines a new paradigm for engineering exotic ordered states beyond conventional supersolids.
@@
To identify different phases, we compute the following observables. The average filling is defined as $\langle n\rangle=\frac{1}{L}\sum_{x}\langle n_{x}\rangle$, and incompressible plateaus in $\langle n\rangle(\mu/V)$ indicate gapped states. The compressibility is $\kappa=\frac{\beta}{L}(\langle N^{2}\rangle-\langle N\rangle^{2})$, with $N$ total number of particles of the system, estimated via density fluctuations, with vanishing $\kappa$ marking insulating behavior. The superfluid density is $\rho_{s}=\langle W^{2}\rangle/(2\beta t)$ [27], where $W$ is the winding number of worldlines. Here, $\rho_{s}>0$ implies global phase coherence.
To characterize spatial ordering we compute the imaginary-time–averaged connected density–density correlator $C(x)=\frac{1}{L}\sum_{i}\big(\langle n_{i}n_{i+x}\rangle-\langle n_{i}\rangle\langle n_{i+x}\rangle\big)-\langle n\rangle$. Its Fourier transform, the static structure factor, $S(q)=\frac{1}{L}\sum_{x}e^{iqx}C(x)$, provides momentum-resolved information on density modulations. In periodic systems $S(q)$ exhibits Bragg peaks at commensurate momenta, while in quasiperiodic systems sharp incommensurate peaks signal quasi–long-range order. In our model these peaks appear at golden–ratio–related wavevectors $q^{*}$, offering a clear signature of quasi-solid or quasi-supersolid phases stabilized by interaction-driven quasiperiodicity. We further cross-check the ordering wavevectors by performing an independent Fourier analysis of the (imaginary-time–averaged) real-space density profile, $\langle n_{x}\rangle\equiv\frac{1}{\beta}\int_{0}^{\beta}d\tau\,\langle n_{x}(\tau)\rangle$. Specifically, we compute $\tilde{n}(q)=\frac{1}{L}\sum_{x}e^{iqx}\big(\langle n_{x}\rangle-\langle n\rangle\big)$. The dominant peaks in $\tilde{n}(q)$ occur at the same incommensurate momenta as those in $S(q)$, confirming that the peaks reflect genuine density modulations rather than artifacts of a particular estimator.
@@
Figure 1: Ground-state phase diagram of the system in the $\mu/V$$t/V$ plane for $L=233$. Two incompressible quasi-solid (qS) lobes emerge near fillings $\langle n\rangle=0.618\approx\alpha$ and $\langle n\rangle=0.382\approx 1-\alpha$, where $\alpha=(1-\sqrt{5})/2$ is the inverse golden ratio. These phases exhibit vanishing compressibility and superfluid density, along with sharp incommensurate peaks in $S(q)$, indicating Fibonacci-induced spatial order. The light blue region marks the quasi-supersolid (qSS) regime, where incommensurate density modulations coexist with finite superfluid density.
Phase diagram– Figure 1 summarizes the ground-state phase diagram of a one-dimensional hard-core Bose system with Fibonacci-modulated quasiperiodic long-range interactions. For system size $L=233$, we map out the phases in the $\mu/V$$t/V$ plane and identify two insulating lobes near fillings $\langle n\rangle\approx\alpha$ and $1-\alpha$, where $\alpha$ is the inverse golden ratio. These lobes exhibit zero compressibility ($\kappa=0$), vanishing superfluid density ($\rho_{s}=0$), and sharp incommensurate Bragg peaks in S(q)S(q)—hallmarks of quasi-solid phases stabilized solely by deterministic incommensurate interactions.
@@
Figure 2: Average filling $\langle n\rangle$ (green circles), superfluid density $\rho_{s}$ (red squares), and structure factor at the dominant incommensurate peak $S(q^{*})$ (purple upward triangles) as functions of chemical potential $\mu/V$ for $L=233$ at $t/V=0.5$. Two prominent incompressible plateaus appear at irrational fillings $\langle n\rangle=0.618$ and $\langle n\rangle=0.382$, corresponding to the inverse golden-ratio $\alpha$ and $1-\alpha$ with $\alpha=(1-\sqrt{5})/2$. These densities are naturally commensurate with the underlying Fibonacci-modulated interaction pattern and host quasi-solid order, characterized by vanishing $\rho_{s}$ and enhanced $S(q^{*})$. Between these plateaus, finite $\rho_{s}$ coexists with nonzero $S(q^{*})$, signaling the emergence of a quasi-supersolid phase stabilized purely by quasiperiodic long-range interactions. Error bars are within symbols if not seen in the figure.
@@
Figure 3: (a) Site-resolved average density profile $\langle n_{x}\rangle-\langle n\rangle$ along a one-dimensional chain of length $L=233$, with total particle number $N=144$, corresponding to filling $\langle n\rangle=0.618\approx\alpha$, where $\alpha=(1-\sqrt{5})/2$ is the inverse golden ratio at $t/V=1/3$. The quasiperiodic long-range interaction follows a Fibonacci modulation pattern, leading to a quasi-solid phase characterized by incommensurate density modulations, vanishing superfluid density, and zero compressibility. (b) Zoom-in of sites 1–54 of the same system. The circle radii are proportional to the local density at each site, emphasizing the incommensurate spatial modulation.
@@
Figure 4: (a) Density–density correlation function $C(x)$ at $t/V=1/3$ for three representative fillings on a system of size $L=233$: $\langle n\rangle=0.618$ (red circles), $\langle n\rangle=0.694$ (blue squares), and $\langle n\rangle=0.780$ (green upward triangles). Red data correspond to the inverse golden ratio $\langle n\rangle=0.618\approx\alpha$, at which quasiperiodic modulation is most pronounced. Fitting curves are superimposed on the data. (b) Corresponding structure factor $S(q)$ for each filling, displaying prominent peaks at incommensurate wavevectors $q_{1}=2.398$ and $q_{2}=3.882$ that reflect the dominant Fourier components of the Fibonacci pattern. Inset: finite-size scaling of the dominant peak height $S(q^{*})$ at $q^{*}=2.398$ for $\langle n\rangle=0.618$ with the fitting function as $S(q^{*})(L)=0.126865+12.3117\,\left(\frac{1}{L}\right)^{1.42065}$, which confirms the persistence of quasi-long-range order in the thermodynamic limit.
@@
To further characterize the nature of the emergent quasi-solid and quasi-supersolid phases, we analyze the time averaged connected density–density correlation function $C(x)=\frac{1}{L}\sum_{i}(\langle n_{i}n_{i+x}\rangle-\langle n_{i}\rangle\langle n_{i+x}\rangle)-\langle n\rangle$ and its Fourier transform $S(q)$ at $t/V=1/3$ for system size $L=233$. Figure 4(a) shows $C(x)$ at three representative fillings: $\langle n\rangle=0.618$, $0.694$, and $0.780$. At $\langle n\rangle=0.618\approx\alpha$, $C(x)$ displays coherent quasiperiodic oscillations over long distances, in stark contrast to the more damped behavior at nearby fillings. We fit $C(x)$ using the function $C(x)=A\cos(qx)e^{-x/\xi}+C_{0}$ for each case. For $\langle n\rangle=0.618$, the best-fit parameters are $A=0.0837$, $q=2.4$, $\xi\approx 185001$, and $C_{0}\approx-2.34\times 10^{-7}$; for $\langle n\rangle=0.694$, the fitting yields $A=0.0334$, $q=2.40001$, $\xi\approx 31615.1$, and $C_{0}\approx-5.34\times 10^{-7}$; while for $\langle n\rangle=0.780$, we obtain $A=0.0122$, $q=2.40001$, $\xi\approx 4962.7$, and $C_{0}\approx-1.21\times 10^{-6}$. The gradual reduction in both the amplitude $A$ and correlation length $\xi$ as the filling deviates from the Fibonacci value does not imply the disappearance of ordered phases. Rather, it reflects a smooth transition from a strongly correlated quasi-solid state at $\langle n\rangle=0.618$—where density modulations dominate and $\rho_{s}=0$—to a quasi-supersolid regime at nearby incommensurate fillings. In this quasi-supersolid regime, the modulated density pattern persists while phase coherence emerges, as evidenced by finite values of superfluid density $\rho_{s}=0.297\pm 0.0008$, compressibility $\kappa=0.323\pm 0.007$ at $\langle n\rangle=0.694$, and $\rho_{s}=0.247\pm 0.0003$, $\kappa=0.384\pm 0.009$ at $\langle n\rangle=0.780$. This evolution highlights the robustness of quasiperiodic interaction–induced ordering: while the strength of density modulation weakens away from the Fibonacci resonance, the system continuously retains quasi-supersolid phase at $t/V=1/3$.
@@
For a system of length $L=F_{k}$, where $F_{k}$ denotes the kk‑th Fibonacci number, the leading Fourier components of the quasiperiodic modulation occur at $q^{*}=2\pi\frac{F_{k-2}}{F_{k}}$ and its mirror $2\pi-q^{*}$. These characteristic momenta arise from the Fourier transform of the Fibonacci characteristic sequence and are a hallmark of its singular‑continuous diffraction spectrum [28, 29, 30]. In our case with $L=233=F_{13}$, the dominant peak appears at $q^{*}=2.398$, corresponding to $2\pi\times 89/233$, confirming that the observed density pattern is intrinsically locked to the inverse golden ratio $\alpha=(1-\sqrt{5})/2$ through its rational approximants $F_{k-2}/F_{k}$. The self‑similar nature of the Fibonacci sequence therefore governs the emergence of quasi‑long‑range spatial order. In addition to these principal peaks, two weaker features appear at $q_{3}=1.483$ and $q_{4}=4.800$. These originate from higher‑order Fourier components of the singular‑continuous spectrum — specifically from secondary harmonics associated with other Fibonacci ratios such as $F_{k-3}/F_{k}$, and reflect the hierarchical, self‑similar structure of the underlying quasiperiodic modulation.
Upon doping away from $\langle n\rangle=\alpha$, the Bragg peaks gradually decreases while superfluid density remains finite, signaling a transition to a quasi-supersolid regime. This demonstrates that quasiperiodic long-range interactions can stabilize both incompressible quasi-solids and nearby quasi-supersolid phases with simultaneous density order and phase coherence.
Discussion and Outlook– Our results demonstrate that deterministic quasiperiodic long-range interactions—without any external potentials, disorder, or soft-core effects—are sufficient to stabilize two exotic many-body phases: the quasi-solid and the quasi-supersolid. These ordered phases emerge solely from the incommensurate structure encoded in the interaction profile and are characterized by the coexistence of spatial density modulation and long-range superfluid coherence. This establishes interaction-induced quasiperiodicity as a new organizing mechanism for stabilizing quasicrystalline quantum phases of matter.
@@
“Interaction-Induced Quasicrystalline Order: Emergence of Quasi-Solid and Quasi-Supersolid Phases”
@@
with total particle number $N=\sum_{x}\langle n_{x}\rangle$. Vanishing $\kappa$ signals incompressibility. We also perform an independent Fourier analysis of the (imaginary-time–averaged) real-space density profile, $\langle n_{x}\rangle\equiv\frac{1}{\beta}\int_{0}^{\beta}d\tau\,\langle n_{x}(\tau)\rangle$. Specifically, we compute $\tilde{n}(q)=\frac{1}{L}\sum_{x}e^{iqx}\big(\langle n_{x}\rangle-\langle n\rangle\big)$.
@@
In particular, for $L=233=F_{13}$, the dominant peaks appear at $q^{*}=2\pi\cdot 89/233=2.398$ and its mirror $2\pi-q^{*}=3.882$, consistent with the main features observed in $S(q)$. This confirms that the emergent density order is intrinsically locked to the inverse golden ratio $\alpha=(1-\sqrt{5})/2$ via its rational approximants $F_{k-2}/F_{k}$.
In addition to these dominant peaks, we also observe subleading features at $q=1.483$ and $4.800$, which correspond to higher-order harmonics at $2\pi\cdot F_{k-3}/F_{k}=2\pi\cdot 55/233$ and its mirror. Together, the presence of both dominant and subleading peaks confirms that the system exhibits quasiperiodic density order governed by multiple competing incommensurate wavevectors. Moreover, the dominant peak positions extracted from $S(q)$ coincide with those obtained from the Fourier spectrum of the imaginary-time–averaged density map $|\tilde{n}(q)|$, confirming that these peaks reflect genuine density modulations rather than artifacts of a particular estimator.
@@
Figure 5: (a) Average filling $\langle n\rangle$, superfluid density $\rho_{s}$, and structure factor $S(q^{*})$ as functions of the chemical potential $\mu/V$ for $t/V=2$ at system size $L=233$. The system is predominantly superfluid, exhibiting finite $\rho_{s}$, while $S(q^{*})$ remains nearly zero across most of the chemical potential range. A weak enhancement of $S(q^{*})$ ($\max S(q^{*})\approx 0.035$) occurs near the inverse golden-ratio fillings $\langle n\rangle=0.382$ and $\langle n\rangle=0.618$, indicating the quasiperiodic order. (b) Corresponding data for $t/V=1/3$. In this regime, well-defined incompressible plateaus appear at $\langle n\rangle=0.382$ and $\langle n\rangle=0.618$, accompanied by enhanced $S(q^{*})$, marking the emergence of stable quasi-solid phases. Between the plateaus, $\rho_{s}$ remains finite while $S(q^{*})$ stays nonzero, signifying a quasi-supersolid region where phase coherence coexists with incommensurate density modulation.
@@
Weak Interaction Regime (t/V=2t/V=2)– Figure S1(a) shows the behavior of $\langle n\rangle$, $\rho_{s}$, and $S(q^{*})$ as functions of $\mu/V$ in the weak-interaction regime. The system is predominantly superfluid: $\rho_{s}$ is finite, and $S(q^{*})$ stays very small over most of the $\mu/V$ range. However, near fillings $\langle n\rangle=0.618\approx\alpha$ and $0.382=1-\alpha$, $S(q^{*})$ develops a shallow maximum (peaking at $\sim 0.035$) and $\langle n\rangle(\mu/V)$ exhibits narrow plateaus. According to our conservative criterion ($S(q^{*})>0.03$), these windows mark the onset of quasi-solid correlations and incipient incompressibility, identifying them as quasi solid regions. Outside these plateaus, $S(q^{*})$ quickly falls below 0.03, indicating that the system remains a compressible superfluid.